OpenStudy (anonymous):

A school captain and 2 vice-captains are to be chosen from a group of 5 boys and 6 girls. What is the probability that all 3 positions will be taken by: at least two girls?

OpenStudy (anonymous):

Here we will use the concept that I was applying that day I guess..

OpenStudy (anonymous):

Firstly find the total number of combinations you can choose to form Captain and Vice captains.. You can do that in : \[\large ^{11}C_3\]

OpenStudy (anonymous):

And note that the word "Atleast " comes here..

OpenStudy (anonymous):

3 positions taken by at least 2 girls has the following 2 cases: 1) 2G and 1B 2) 3G and 0B no. of ways for (1) is 6C2 * 5C1 no. of ways for (2) is 6C3 * 5C0 so the total no. of ways for at least 2 girls is (6C2 * 5C1) + (6C3 * 5C0) for the probability, divide that exp by the total no. of ways for selecting the captains = 11C3

OpenStudy (anonymous):

Atleast means you can have 2 girls out of 3 or you can have 3 out of 3 also.. SO: \[^6C_2 \times ^5C_1 + ^6C_3\]

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

sorry, explain please :(

OpenStudy (anonymous):

which part is confusing u?

OpenStudy (anonymous):

I am making the combinations so check this: 2 girls And 1 boy Or 3 girls AND 0 boy.. \(^6C_2\) \(\times \) \(^5C_1\) + \(^6C_3\) \(\times\) \(^5C_0\) 2 girls out of 6 you can choose in \(^6C_2\) 1 boy you can select in out of 5 : \(^5C_1\) 3 girls out of 6 you can have in : \(^6C_3\) 0 boys out of 5 you can have in : \(^5C_0\)

OpenStudy (anonymous):

ok got it, thanks.

OpenStudy (anonymous):

Sure no ??

OpenStudy (anonymous):

ya, sure.

OpenStudy (anonymous):

Solve it and try to find out the answer..

OpenStudy (anonymous):

yep, got it.

OpenStudy (anonymous):

Well Done..